# Personalized Federated Learning with Heat-Kernel Enhanced Tensorized Multi-View Clustering

### v1

@@ 2.3 Heat Kernel Methods and Quantum-Inspired Learning
The heat kernel, defined as the fundamental solution to the heat equation, has found extensive applications in machine learning through kernel methods and manifold learning. Ham et al. [28] introduced diffusion maps that leverage heat kernel properties for dimensionality reduction, while Coifman and Lafon [29] established theoretical foundations for diffusion-based data analysis.
In clustering applications, heat kernel methods have been used to capture local neighborhood structure and improve cluster boundary detection. Zelnik-Manor and Perona [30] developed self-tuning spectral clustering using local heat kernel scales, while von Luxburg [31] provided comprehensive analysis of heat kernel properties in graph-based clustering.
Quantum-inspired machine learning has gained attention for its ability to leverage quantum mechanical principles in classical algorithms. Quantum kernels, introduced by Havlíček et al. [32], enable the computation of quantum feature maps for enhanced pattern recognition. The connection between quantum field theory and machine learning has been explored through path integral formulations [33] and quantum neural networks [34].
@@ 2.4 Federated Multi-View Learning and Privacy Preservation
Recent work on federated multi-view learning includes approaches by Chen et al. [35] for federated multi-view representation learning and Liu et al. [36] for privacy-preserving multi-view clustering. However, these methods do not leverage tensor decomposition for efficient representation and communication.
@@
So, we will discuss the tensorization and activation of Eq. 6 in federated environments, addressing these open challenges through our proposed heat-kernel enhanced tensorized framework in the next sections.
@@ 2.5 Tensor Decomposition Methods
Canonical Polyadic Decomposition (CPD), also known as CANDECOMP/PARAFAC, represents the foundational tensor decomposition method. Introduced independently by Carroll and Chang [37] and Harshman [38], CPD decomposes a tensor into a sum of rank-one tensors, providing a unique factorization under mild conditions. The uniqueness property of CPD makes it particularly valuable for component analysis and factor identification in multi-way data.
Tucker decomposition, proposed by Tucker [39], provides a more flexible alternative through its core tensor and factor matrices structure. Unlike CPD, Tucker decomposition allows different ranks along each mode, making it suitable for data with varying complexity across dimensions. The Higher-Order SVD (HOSVD) algorithm by De Lathauwer et al. [40] established efficient computational procedures for Tucker decomposition, enabling its application to large-scale problems.
Recent advances in tensor decomposition have focused on addressing practical challenges such as missing data, noise robustness, and computational scalability. Tensor completion methods, including work by Liu et al. [41] and Yuan and Zhang [42], enable handling of incomplete tensors through low-rank assumptions. Robust tensor decomposition techniques [43] provide resilience against outliers and corrupted entries.
Non-negative tensor factorization (NTF) has gained attention for applications requiring interpretable, parts-based representations. Lee and Seung’s pioneering work [44] on non-negative matrix factorization was extended to tensors by Shashua and Hazan [45], enabling decompositions with non-negativity constraints that often yield more interpretable factors.
The integration of tensor decomposition with machine learning has led to tensorized neural networks [46], tensor-based kernel methods [47], and multi-task learning frameworks [48]. These approaches leverage tensor structure to reduce model complexity while maintaining representational power.
In the context of clustering, tensor-based methods have shown particular promise for multi-view and multi-way data analysis. Zhou et al. [49] developed tensor-based multiway arrays for bid data sets, while Benjamin et al. [50] proposed tensorized possibilistic c-means for high-dimensional clustering. However, these methods are not designed for federated environments and lack the privacy-preserving mechanisms required for distributed deployment.
@@
Tensor Operations and Decomposition Preliminaries: The elements of $N$-way arrays can be reordered into matrices or vectors through matricization and vectorization operations. The matricization is reordering tensor’ elements into a matrix, while vectorization is reordering tensor’ elements into a vector. Canonical polyadic decomposition (CPD) or parallel factors (PARAFAC) [51, 37], Tucker decomposition (TD), and non-negative tensor factorization (NTF) are well-known approaches to reorder a tensor into a matrix or vector. PARAFAC or CPD holds multilinearity assumptions such as handle a sum of rank-one tensors by targeting the same lengths. In this sense, the original PARAFAC decomposition can reveal uncorrelated (orthogonal) and correlated (oblique) factors in system-variation data matrices, but failed to describe correlated factors when applied to the cross-product matrices computed from one data [52, 53]. PARAFAC2, on the other hand has the advantageous to handle irregular tensors by relaxing the canonical polyadic (CP) model [52]. PARAFAC2 enables the three-mode model, often implemented into subspace clustering algorithms taking advantageous of correlations between pair of factors. The PARAFAC2 normalize the generalization of PARAFAC to deal both with uncorrelated and correlated factors.
@@ 2.5.1 The Basic Background of Tensors
This subsection provides some definitions and concepts of tensor algebra, most of overviews and notations are borrowed from Kolda et al. [54].
@@
where the $(g_{1},g_{2},\ldots,g_{n})$-th elements of the tensor $\mathcal{A}\in\mathbb{R}^{G_{1}\times G_{2}\times\ldots\times G_{n}\times\ldots\times G_{N}}$ is containing $(\mathcal{A})_{g_{1},g_{2},\ldots,g_{n}}$ or $a_{g_{1},g_{2},\ldots,g_{n}}$ for indices $g=\{1,2,\ldots,G_{1}\}$, $g=\{1,2,\ldots,G_{2}\}$, and $g_{n}=\{1,2,\ldots,G_{N}\}$.
@@ 2.5.2 The Higher Order Singular Value Decomposition (HOSVD)
Tensor factorization had been implemented as a decomposition-based approach to project a large amount of data into a small number of data considering collinearity of its heterogeneous [55]. CANDECOMP/PARAFAC is one of the most existing tensor decomposition techniques to improve data clustering processes [54, 56]. The HOSVD is a multidimensional extension to the matrix singular value decomposition (SVD). To better explain the higher-order SVD, we will borrow some concept of Wang and Song [55] with a two-mode of tensor order $N$.
@@ 2.6 Summary and Research Gaps
The review of existing literature reveals several important research gaps that motivate our proposed approach:
Limited Integration: While federated learning, multi-view clustering, and tensor decomposition have been extensively studied independently, their integration remains largely unexplored. Existing federated clustering methods do not leverage multi-view structure, and multi-view clustering approaches are not designed for federated environments.
Communication Inefficiency: Current federated multi-view methods do not exploit tensor structure for communication compression. The high-dimensional nature of multi-view data exacerbates communication overhead in federated settings, requiring novel tensor-aware compression schemes.
Privacy Limitations: Existing privacy-preserving mechanisms are not optimized for multi-view tensor data, potentially leading to information leakage across views or reduced clustering utility. New differential privacy mechanisms tailored to tensor structures are needed.
Lack of Personalization: Most federated clustering approaches assume homogeneous data distributions across clients. In reality, clients often have heterogeneous data characteristics requiring personalized clustering solutions that adapt to local data patterns while benefiting from global knowledge.
Theoretical Understanding: The convergence properties and privacy guarantees of federated multi-view tensor clustering are not well understood. Theoretical analysis is needed to provide convergence guarantees and characterize the privacy-utility trade-offs.
Our work addresses these gaps by proposing a novel heat-kernel enhanced tensorized framework that integrates advanced tensor decomposition with personalized federated learning, providing both theoretical guarantees and practical advantages in communication efficiency and clustering performance.
3 Problem Formulation
@@ 3.3 Personalized Federated Learning
$J_{E-FKMVC}(V_{[\ell]},U_{[\ell]}^{*},A_{[\ell]}^{h})=\sum_{\ell=1}^{M}\sum_{h=1}^{s(\ell)}v_{[\ell]h}^{\alpha}\sum_{i=1}^{n(\ell)}\sum_{k=1}^{c(\ell)}(\mu_{[\ell]ik}^{*})^{m}\text{~FKED}(x_{[\ell]ij}^{h},a_{[\ell]kj}^{h})$ (22)
@@
$\begin{split}\tilde{J}_{E-FKMVC}^{\ell}&=J_{E-FKMVC}(V_{[\ell]},U_{[\ell]}^{*},A_{[\ell]}^{h})\\ &\quad+\sum_{i=1}^{n(\ell)}{\lambda_{1i}}\left(\sum_{k=1}^{c(\ell)}{\mu_{[\ell]ik}^{*}-1}\right)+{\lambda_{2}}\left(\sum_{h=1}^{s(\ell)}{v_{[\ell]h}-1}\right)\end{split}$ (25)
@@
$\frac{\partial\tilde{J}_{E-FKMVC}^{\ell}}{\partial\mu_{[\ell]ik}^{*}}$ $=\frac{\partial}{\partial\mu_{[\ell]ik}^{*}}\left[\sum_{h=1}^{s(\ell)}v_{[\ell]h}^{\alpha}(\mu_{[\ell]ik}^{*})^{m}\text{FKED}(x_{[\ell]i}^{h},a_{[\ell]k}^{h})\right]+\lambda_{1i}$ (29)
@@
$\frac{\partial J_{E-FKMVC}^{\ell}}{\partial a_{[\ell]kj}^{h}}$ $=\frac{\partial}{\partial a_{[\ell]kj}^{h}}\left[v_{[\ell]h}^{\alpha}\sum_{i=1}^{n(\ell)}(\mu_{[\ell]ik}^{*})^{m}\text{FKED}(x_{[\ell]i}^{h},a_{[\ell]k}^{h})\right]$ (37)
@@
$\frac{\partial\tilde{J}_{E-FKMVC}^{\ell}}{\partial v_{[\ell]h}}$ $=\frac{\partial}{\partial v_{[\ell]h}}\left[\sum_{h^{\prime}=1}^{s(\ell)}v_{[\ell]h^{\prime}}^{\alpha}\sum_{i=1}^{n(\ell)}\sum_{k=1}^{c(\ell)}(\mu_{[\ell]ik}^{*})^{m}\text{FKED}(x_{[\ell]i}^{h^{\prime}},a_{[\ell]k}^{h^{\prime}})\right]+\lambda_{2}$ (49)
@@
    Stationarity: $\nabla J_{E-FKMVC}^{\ell}+\sum_{i=1}^{n(\ell)}\lambda_{1i}\nabla g_{i}+\lambda_{2}\nabla h=0$, where $g_{i}(\mu)=\sum_{k=1}^{c(\ell)}\mu_{[\ell]ik}^{*}-1$ and $h(v)=\sum_{h=1}^{s(\ell)}v_{[\ell]h}-1$.
@@
  • Boundedness and Continuity: The objective function $J_{E-FKMVC}^{\ell}$ is continuous and bounded below. This follows from the fact that FKED values are bounded in $[0,1]$ by construction (Eq. 4), ensuring the weighted sum in the objective function remains finite. The exponential decay in the heat kernel guarantees that distances approach finite limits as data points move infinitely far apart.
@@
3.4 Tensorized Federated Objective
Based on Eq. 22, we propose the global objective combines local tensorized mullti-view clustering with federated coordination (T-MVCF):
$\min_{\{\mathcal{U}_{[\ell]},\mathcal{A}_{[\ell]},\mathcal{V}_{[\ell]}\},\{\mathcal{A}_{g},\mathcal{V}_{g}\}}\sum_{\ell=1}^{M}\left[\omega_{[\ell]}\mathcal{L}_{tensor}^{(\ell)}(\mathcal{X}_{[\ell]},\mathcal{U}_{[\ell]},\mathcal{A}_{[\ell]},\mathcal{V}_{[\ell]})+\gamma\mathcal{R}_{tensor}(\mathcal{A}_{[\ell]},\mathcal{A}_{g})+\eta\mathcal{R}_{view}(\mathcal{V}_{[\ell]},\mathcal{V}_{g})\right]$ (57)
where:
  • $\mathcal{L}_{tensor}^{(\ell)}$ is the local tensorized heat-kernel enhanced clustering loss
  • $\mathcal{U}_{[\ell]}$ represents the membership tensor for client $\ell$
  • $\mathcal{A}_{[\ell]}$ represents the tensor of cluster centers for client $\ell$
  • $\mathcal{V}_{[\ell]}$ represents the view weight tensor for client $\ell$
  • $\mathcal{R}_{tensor}$ and $\mathcal{R}_{view}$ are tensorized regularization terms ensuring global coordination
3.5 Tensor Decomposition Framework
To handle the high-dimensional tensor structures efficiently, we employ tensor decomposition techniques. For client $\ell$, we decompose the cluster center tensor using Tucker decomposition:
$\mathcal{A}_{[\ell]}\approx\mathcal{G}_{[\ell]}\times_{1}\mathbf{U}_{[\ell]}^{(1)}\times_{2}\mathbf{U}_{[\ell]}^{(2)}\times_{3}\mathbf{U}_{[\ell]}^{(3)}$ (58)
where $\mathcal{G}_{[\ell]}$ is the core tensor and $\mathbf{U}_{[\ell]}^{(z)}$ are the factor matrices for mode $z$. This decomposition enables:
  1. 1.
    Dimensionality reduction through low-rank approximation
  2. 2.
    Efficient storage and communication of tensor parameters
  3. 3.
    Capture of multilinear relationships across different data modes

### v2

@@ 3.3 Heat Kernel Methods and Quantum-Inspired Learning
The heat kernel, defined as the fundamental solution to the heat equation, has found extensive applications in machine learning through kernel methods and manifold learning. Ham et al. [ham2004kernel] introduced diffusion maps that leverage heat kernel properties for dimensionality reduction, while Coifman and Lafon [coifman2006diffusion] established theoretical foundations for diffusion-based data analysis.
In clustering applications, heat kernel methods have been used to capture local neighborhood structure and improve cluster boundary detection. Zelnik-Manor and Perona [zelnik2004self] developed self-tuning spectral clustering using local heat kernel scales, while von Luxburg [von2007tutorial] provided comprehensive analysis of heat kernel properties in graph-based clustering.
Quantum-inspired machine learning has gained attention for its ability to leverage quantum mechanical principles in classical algorithms. Quantum kernels, introduced by Havlíček et al. [havlivcek2019supervised], enable the computation of quantum feature maps for enhanced pattern recognition. The connection between quantum field theory and machine learning has been explored through path integral formulations [carleo2017solving] and quantum neural networks [biamonte2017quantum].
@@ 3.4 Federated Multi-View Learning and Privacy Preservation
Recent work on federated multi-view learning includes approaches by Chen et al. [chen2022federated] for federated multi-view representation learning and Liu et al. [liu2023federated] for privacy-preserving multi-view clustering. However, these methods do not leverage tensor decomposition for efficient representation and communication.
@@
The subsequent discourse will encompass the tensorization and activation of equation Eq. 6. In the following sections, the open challenges will be addressed by means of a proposed heat-kernel-enhanced tensorized framework in federated environments.
@@ 3.5 Tensor Decomposition Methods
Canonical Polyadic Decomposition (CPD), also known as CANDECOMP/PARAFAC, represents the foundational tensor decomposition method. Introduced independently by Carroll and Chang [carroll1970analysis] and Harshman [harshman1970foundations], CPD decomposes a tensor into a sum of rank-one tensors, providing a unique factorization under mild conditions. The uniqueness property of CPD makes it particularly valuable for component analysis and factor identification in multi-way data.
Tucker decomposition, proposed by Tucker [tucker1966some], provides a more flexible alternative through its core tensor and factor matrices structure. Unlike CPD, Tucker decomposition allows different ranks along each mode, making it suitable for data with varying complexity across dimensions. The Higher-Order SVD (HOSVD) algorithm by De Lathauwer et al. [de2000multilinear] established efficient computational procedures for Tucker decomposition, enabling its application to large-scale problems.
Recent advances in tensor decomposition have focused on addressing practical challenges such as missing data, noise robustness, and computational scalability. Tensor completion methods, including work by Liu et al. [liu2012tensor] and Yuan and Zhang [yuan2016tensor], enable handling of incomplete tensors through low-rank assumptions. Robust tensor decomposition techniques [huang2014robust] provide resilience against outliers and corrupted entries.
Non-negative tensor factorization (NTF) has gained attention for applications requiring interpretable, parts-based representations. Lee and Seung’s pioneering work [lee2000algorithms] on non-negative matrix factorization was extended to tensors by Shashua and Hazan [shashua2005non], enabling decompositions with non-negativity constraints that often yield more interpretable factors.
The integration of tensor decomposition with machine learning has led to tensorized neural networks [novikov2015tensorizing], tensor-based kernel methods [signoretto2011kernel], and multi-task learning frameworks [wimalawarne2014multitask]. These approaches leverage tensor structure to reduce model complexity while maintaining representational power.
In the context of clustering, tensor-based methods have shown particular promise for multi-view and multi-way data analysis. Zhou et al. [cichocki2015tensor] developed tensor-based multiway arrays for bid data sets, while Benjamin et al. [benjamin2024tensor] proposed tensorized possibilistic c-means for high-dimensional clustering. However, these methods are not designed for federated environments and lack the privacy-preserving mechanisms required for distributed deployment.
@@
Tensor Operations and Decomposition Preliminaries: The elements of $N$-way arrays can be reordered into matrices or vectors through matricization and vectorization operations. The matricization is reordering tensor’ elements into a matrix, while vectorization is reordering tensor’ elements into a vector. Canonical polyadic decomposition (CPD) or parallel factors (PARAFAC) [harshman1994parafac, carroll1970analysis], Tucker decomposition (TD), and non-negative tensor factorization (NTF) are well-known approaches to reorder a tensor into a matrix or vector. PARAFAC or CPD holds multilinearity assumptions such as handle a sum of rank-one tensors by targeting the same lengths. In this sense, the original PARAFAC decomposition can reveal uncorrelated (orthogonal) and correlated (oblique) factors in system-variation data matrices, but failed to describe correlated factors when applied to the cross-product matrices computed from one data [harshman1972parafac2, kiers1999parafac2]. PARAFAC2, on the other hand has the advantageous to handle irregular tensors by relaxing the canonical polyadic (CP) model [harshman1972parafac2]. PARAFAC2 enables the three-mode model, often implemented into subspace clustering algorithms taking advantageous of correlations between pair of factors. The PARAFAC2 normalize the generalization of PARAFAC to deal both with uncorrelated and correlated factors.
@@ 3.5.1 The Basic Background of Tensors
This subsection provides some definitions and concepts of tensor algebra, most of overviews and notations are borrowed from Kolda et al. [kolda2009tensor].
@@
where the $(g_{1},g_{2},\ldots,g_{n})$-th elements of the tensor $\mathcal{A}\in\mathbb{R}^{G_{1}\times G_{2}\times\ldots\times G_{n}\times\ldots\times G_{N}}$ is containing $(\mathcal{A})_{g_{1},g_{2},\ldots,g_{n}}$ or $a_{g_{1},g_{2},\ldots,g_{n}}$ for indices $g_{1}=\{1,2,\ldots,G_{1}\}$, $g_{2}=\{1,2,\ldots,G_{2}\}$, and $g_{n}=\{1,2,\ldots,G_{N}\}$.
@@ 3.5.2 The Higher Order Singular Value Decomposition (HOSVD)
Tensor factorization is a method for reducing a large volume of data to a smaller set of data points. This is achieved by leveraging the collinearity of its heterogeneous data [wang2017tensor]. CANDECOMP/PARAFAC is a technique for decomposing data into its constituent parts. It is one of the most widely used tensor decomposition techniques, as referenced in the works of Kolda et al. (2009) [kolda2009tensor]. and Kiers (2000) [kiers2000towards]. Conversely, the HOSVD represents an extension of the matrix singular value decomposition (SVD). In order to provide a more thorough explanation of the higher-order SVD (HOSVD), this paper will draw upon certain concepts from Wang and Song’s work on two-mode tensor orders $N$ [wang2017tensor].
@@ 3.6 Summary and Research Gaps
A review of extant literature reveals several significant research gaps that motivate the proposed approach, including the limited integration of research findings. While federated learning, multi-view clustering, and tensor decomposition have been the focus of extensive research study, their integration remains largely unexplored. Existing federated clustering methods do not leverage multi-view structure, and multi-view clustering approaches are not designed for federated environments. In the context of communication inefficiency, contemporary federated multi-view methods do not leverage tensor structure for the purpose of communication compression. The high-dimensional nature of multi-view data exacerbates communication overhead in federated settings, requiring novel tensor-aware compression schemes.
Existing privacy-preserving mechanisms are not optimized for multi-view tensor data, potentially leading to information leakage across views or reduced clustering utility. The development of novel differential privacy mechanisms tailored to tensor structures is imperative. A further issue is the absence of personalization. The prevailing assumption underlying most federated clustering approaches is that data distributions across clients are homogeneous. In practice, clients frequently possess heterogeneous data characteristics, necessitating customized clustering solutions that adapt to local data patterns while leveraging global knowledge. As indicated in the theoretical understanding, the convergence properties and privacy guarantees of federated multi-view tensor clustering remain to be fully elucidated. A theoretical analysis is necessary to provide convergence guarantees and characterize the privacy-utility trade-offs.
The proposed frameworks address these limitations by integrating advanced tensor decomposition with personalized federated learning, leveraging a heat-kernel-enhanced tensorized approach. The aforementioned framework provides both theoretical guarantees and practical advantages.
4 Proposed Methodology
This section presents the proposed heat-kernel enhanced tensorized framework for personalized federated multi-view clustering. The initial step involves formulating the problem setting and introducing the federated kernel Euclidean distance metric. In the following section, the calculation of federated heat-kernel coefficients adapted to the multi-view tensor data structure will be delineated. Finally, the overall algorithmic framework must be described, including local updates, global aggregation, and privacy-preserving mechanisms.
4.1 Problem Formulation
@@ 4.3 Federated Heat-Kernel Coefficients
5 The Enhanced Federated Kernel Multi-View Clustering
In this section, the proposed Enhanced Federated Kernel Multi-View Clustering (E-FKMVC) algorithm will be presented. The E-FKMVC algorithm is a refinement of the traditional multi-view clustering framework that operates within a federated learning environment. This refinement enables multiple clients to collaboratively cluster their local multi-view datasets while preserving data privacy.
@@ 5.1 Personalized Federated Learning
$J_{E-FKMVC}^{(\ell)}=J_{E-FKMVC}(V_{[\ell]},U_{[\ell]}^{*},A_{[\ell]}^{h})=\sum_{\ell=1}^{M}\sum_{h=1}^{s(\ell)}v_{[\ell]h}^{\alpha}\sum_{i=1}^{n(\ell)}\sum_{k=1}^{c(\ell)}(\mu_{[\ell]ik}^{*})^{m}\text{~FKED}(x_{[\ell]ij}^{h},a_{[\ell]kj}^{h})$ (22)
@@
$\begin{split}\tilde{J}_{E-FKMVC}^{(\ell)}&=J_{E-FKMVC}(V_{[\ell]},U_{[\ell]}^{*},A_{[\ell]}^{h})\\ &\quad+\sum_{i=1}^{n(\ell)}{\lambda_{1i}}\left(\sum_{k=1}^{c(\ell)}{\mu_{[\ell]ik}^{*}-1}\right)+{\lambda_{2}}\left(\sum_{h=1}^{s(\ell)}{v_{[\ell]h}-1}\right)\end{split}$ (25)
@@
$\frac{\partial\tilde{J}_{E-FKMVC}^{(\ell)}}{\partial\mu_{[\ell]ik}^{*}}$ $=\frac{\partial}{\partial\mu_{[\ell]ik}^{*}}\left[\sum_{h=1}^{s(\ell)}v_{[\ell]h}^{\alpha}(\mu_{[\ell]ik}^{*})^{m}\text{FKED}(x_{[\ell]i}^{h},a_{[\ell]k}^{h})\right]+\lambda_{1i}$ (29)
@@
$\frac{\partial J_{E-FKMVC}^{(\ell)}}{\partial a_{[\ell]kj}^{h}}$ $=\frac{\partial}{\partial a_{[\ell]kj}^{h}}\left[v_{[\ell]h}^{\alpha}\sum_{i=1}^{n(\ell)}(\mu_{[\ell]ik}^{*})^{m}\text{FKED}(x_{[\ell]i}^{h},a_{[\ell]k}^{h})\right]$ (37)
@@
$\frac{\partial\tilde{J}_{E-FKMVC}^{(\ell)}}{\partial v_{[\ell]h}}$ $=\frac{\partial}{\partial v_{[\ell]h}}\left[\sum_{h^{\prime}=1}^{s(\ell)}v_{[\ell]h^{\prime}}^{\alpha}\sum_{i=1}^{n(\ell)}\sum_{k=1}^{c(\ell)}(\mu_{[\ell]ik}^{*})^{m}\text{FKED}(x_{[\ell]i}^{h^{\prime}},a_{[\ell]k}^{h^{\prime}})\right]+\lambda_{2}$ (49)
@@
    Stationarity: $\nabla J_{E-FKMVC}^{(\ell)}+\sum_{i=1}^{n(\ell)}\lambda_{1i}\nabla g_{i}+\lambda_{2}\nabla h=0$, where $g_{i}(\mu)=\sum_{k=1}^{c(\ell)}\mu_{[\ell]ik}^{*}-1$ and $h(v)=\sum_{h=1}^{s(\ell)}v_{[\ell]h}-1$.
@@
  • Boundedness and Continuity: The objective function $J_{E-FKMVC}^{(\ell)}$ is continuous and bounded below. This follows from the fact that FKED values are bounded in $[0,1]$ by construction (Eq. 4), ensuring the weighted sum in the objective function remains finite. The exponential decay in the heat kernel guarantees that distances approach finite limits as data points move infinitely far apart.
@@
Algorithm 1 summarizes the steps of the proposed E-FKMVC algorithm. This algorithm iteratively updates the membership matrices, cluster centers, and view weights for each client until convergence is achieved.
Algorithm 1 Enhanced Federated Kernel Multi-View Clustering (E-FKMVC)
0: Multi-view data $X_{[\ell]}=\{X_{[\ell]}^{h}\}_{h=1}^{s(\ell)}$ for clients $\ell=1,\ldots,M$, number of clusters $c(\ell)$, fuzziness parameter $m>1$, view weight exponent $\alpha>1$, heat-kernel parameters $\delta_{[\ell]ij}^{h}$
0: Membership matrices $U_{[\ell]}^{*}$, cluster centers $A_{[\ell]}^{h}$, view weights $V_{[\ell]}$
1:Initialization: Randomly initialize $U_{[\ell]}^{*}$, $A_{[\ell]}^{h}$, and $V_{[\ell]}$ for all clients $\ell$
2:repeat
3:  for each client $\ell=1$ to $M$ do
4:   Update membership matrix $U_{[\ell]}^{*}$ using Eq. 26
5:   Update cluster centers $A_{[\ell]}^{h}$ using Eq. 27
6:   Update view weights $V_{[\ell]}$ using Eq. 28
7:  end for
8:until convergence criteria met (e.g., changes in $U_{[\ell]}^{*}$, $A_{[\ell]}^{h}$, and $V_{[\ell]}$ below a threshold)